Circular No. 5736

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J in NGC 3031 Spectrograms obtained with the 1.93-m Observatoire de Haute Pro-

vence telescope by Y. Andrillat on Mar. 30.95 UT showed no feature in the region of the near-infrared Ca II triplet. On Mar. 31.00 Hα was detected in emission with an equivalent width of 0.03 nm and a FWHM of 190 km/s.

CCD photometry by P. Prugniel at the 0.8-m Observatoire de Haute Provence telescope on Mar. 31.00 UT gave B = 11.05, B − R = +0.3.

J.-M. Perelmuter, Observatoire du Mont Mégantic, Université de Montréal, notes the presence on Canada-France-Hawaii Telescope 1980 plates and Kitt Peak National Observatory 1990 CCD images of an object within 0′′.2 of the position of SN 1993J quoted by Hartwick et al. on IAUC 5731. The object appears stellar on images with FWHM as small as 1′′.0 and had V = 20.0, B − V = +1.1, V − R = +0.7 (uncertainty 0.1 mag) in Mar. 1990.

Visual magnitude estimates: Mar. 30.94 UT, 10.5 (J. D. Shanklin, Cambridge, England); 31.1, 10 (R. Kohl, Lakewood, NY; independent discovery); 31.13, 10.5 (F. Cianciolo, R. Gearhart and A. Dosaj, Austin, TX).

SUPERNOVA 1993K IN NGC 2223 M. Della Valle, European Southern Observatory; and J. M. Alcalá,

Landessternwarte, Heidelberg, report: “Analysis of a spectrogram (range 480-710 nm, resolution about 0.5 nm) obtained with the 1.52-m telescope at La Silla on Mar. 31.0 UT shows a featureless blue continuum, indicating a possible type II supernova caught during an early stage.”

G. Massone, Osservatorio di Torino, reports the following magnitudes, obtained on Mar. 31.1 UT with the 0.9-m Dutch telescope at La Silla: V = 15.3, B − V = +0.1, V − R = +0.2.

P. C. Schmidtke, Arizona State University, provides further BVRI photometry with the 0.9-m telescope at the Cerro Tololo Interamerican Observatory. Quick-look estimates, starting on Mar. 31.12 UT, gave V = 15.3, B − V = +0.2, V − R = +0.2 (uncertainty 0.1 mag).

1993 March 31 Brian G. Marsden